The density of aluminum at 298 K is 2.70 g/mL. A sheet of aluminum (Al) foil has a total area of 1.060 ft2 and a mass of 4.254 g. What is the thickness of the foil in micrometers?
1
Expert's answer
2020-10-04T09:14:04-0400


VOLUME= MASS÷DENSITY

        =4.254÷2.70

        =1.57ml

convert volume to cm³

1ml=1cm³      1.57ml=1.57cm³

Convert area to cm²

1ft²=929.03cm²        1.06ft² in cm² will be 929.03×1.06= 987.77cm²

Thickness = VOLUME÷AREA

          =1.5756÷987.77 

          =0.0015994cm.

Convert thickness to micrometers.

1cm =1000 micrometers.

0.0015994cm in micrometers will be 

       0.0015994 ×1000=1.5999micrometers